If A~ = 2yˆi - zˆj + 3xkˆ

(i). Find the unit vectors eˆr, eˆθ and eˆz of a cylindrical coordinates in termsof ˆi, ˆj and kˆ.

(ii). Solve for ˆi, ˆj and kˆ in terms of eˆr, eˆθ and eˆz

(iii). Represent the vector A~ in cylindrical coordinates and determine Ar, Aθ

and Az





1
Expert's answer
2021-08-09T16:07:36-0400

Solutiom;

a)

The position vector is given by;

"\vec{r}=x\hat{i}+y\hat{j}+z\hat{k}"

In cylindrical coordinates;

"\vec{r}=rcos\theta\hat{i}+rsin\theta\hat{j}+z\hat{k}"

Hence,we find ;

"\hat{e}_r=\frac{\frac{d\vec{r}}{dr}}{|\frac{d\vec{r}}{dr}|}" "=\frac{cos\theta\hat{i}+sin\theta\hat{j}}{\sqrt{cos^2\theta+sin^2\theta}}" "\hat{e}_r=cos\theta\hat{i}+sin\theta{j}"

"\hat{e}_{\theta}=\frac{\frac{d\vec{r}}{d\theta}}{|\frac{d\vec{r}}{d\theta}|}" "=\frac{-rsin\theta\hat{i}+rcos\theta\hat{j}}{\sqrt {r^2cos^2\theta+r^2sin^2\theta}}"

"\hat{e}_{\theta}=-sin\theta\hat{i}+cos\theta\hat{j}"

"\hat{e}_z=\frac{\frac{d\vec{r}}{dz}}{|\frac{d\vec{r}}{dz}|}" "=\frac{\hat{k}}{1}=\hat{k}"

b)

As calculated;

"\hat{e}_r=cos\theta\hat{i}+sin\theta\hat{j}" ....(i)

"\hat{e}_{\theta}=-sin\theta\hat{i}+cos\theta\hat{j}" ...(ii)

"\hat{e}_z=\hat{k}"

Multiply (i) by "cos\theta" and (ii) by "sin\theta" and subtract;

"cos\theta\hat{e}_r=cos^2\theta\hat{i}+cos\theta sin\theta\hat{j}"

"-sin\theta\hat{e}_\theta=-sin^2\theta\hat{i}+cos\theta sin\theta\hat{j}"

"\overline{cos\theta\hat{e}_r-sin\theta\hat{e}_{\theta}=(cos^2\theta+sin^2\theta)\hat{i}}"

Hence,

"\hat{i}=cos\theta\hat{e}_r-sin\theta\hat e_{\theta}"

Multiply (i) with "sin\theta" and (ii) with "cos\theta" and add;

"sin\theta\hat e_r=sin\theta cos\theta\hat i+sin^2\theta\hat j"

"+cos\theta\hat e_{\theta}=-sin\theta cos\theta\hat i+cos^2\theta\hat j"

"\overline{sin\theta \hat e_r+cos\theta\hat e_{\theta}=(sin^2\theta+cos^2\theta)\hat j}"

Hence,

"\hat j=sin\theta\hat e_r+cos\theta\hat e_{\theta}"

From (iii);

"\hat k=\hat e_z"

c)

Given;

"\vec{A}=2y\hat i-z\hat j+3x\hat k"

In cylindrical coordinates;

"\vec{A}=2(rsin\theta)(cos\theta\hat e_r-sin\theta\hat e_{\theta})-z(sin\theta\hat e_r+cos\theta\hat e_{\theta})+3(rcos\theta)\hat e_z"

By distribution;

"\vec{A}=(2rsin\theta cos\theta-zsin\theta)\hat e_r-(2rsin^2\theta+zcos\theta)\hat e_{\theta}+3rcos\theta\hat e_z"

From which;

"A_r=2rcos\theta sin\theta-zsin\theta"

"A_{\theta}=-2rsin^2\theta-zcos\theta"

"A_z=3rcos\theta"